OpenStudy (anonymous):

a rectangle is 8 inches wider than half the height. the height is h, so the width is 1/2h+8. if h is 36, what is 1/2h+8?

OpenStudy (anonymous):

\[\frac{1}{2} \times 36+8=18+8=26\]